# DRL-ORA: Distributional Reinforcement Learning with Online Risk Adaption

### v4

@@ Abstract
One of the main challenges in reinforcement learning (RL) is that the agent has to make decisions that would influence the future performance without having complete knowledge of the environment. Dynamically adjusting the level of epistemic risk during the learning process can help to achieve reliable policies in safety-critical settings with better efficiency. In this work, we propose a new framework, Distributional RL with Online Risk Adaptation (DRL-ORA). This framework quantifies both epistemic and implicit aleatory uncertainties in a unified manner and dynamically adjusts the epistemic risk levels by solving a total variation minimization problem online. The framework unifies the existing variants of risk adaption approaches and offers better explainability and flexibility. The selection of risk levels is performed efficiently via a grid search using a Follow-The-Leader-type algorithm, where the offline oracle also corresponds to a “satisficing measure” under a specially modified loss function. We show that DRL-ORA outperforms existing methods that rely on fixed risk levels or manually designed risk level adaptation in multiple classes of tasks.
Introduction
@@
Reinforcement learning (RL) algorithms have shown great success in games and simulated environments (Mnih et al. 2015), and have drawn interest for real-world and industrial applications (Mahmood et al. 2018; Zhang et al. 2020). Taking a step further, distributional reinforcement learning (DRL) (Bellemare, Dabney, and Munos 2017) considers the intrinsic distribution of the returns rather than their expectation, as considered in standard RL settings. This approach allows agents to quantify the risk of their policies and enables the development of risk-aware policies by considering the distributional information of their performance (Dabney et al. 2018b). Quantifying both aleatory and epistemic uncertainties of RL crucial for developing risk-aware policies. The former expresses the inherent randomness in the problem, and the latter represents a lack of knowledge about the environment (Mavrin et al. 2019; Clements et al. 2019; Charpentier et al. 2022). However, most existing research has focused on risk-aware RL with a fixed risk level, such as investigating risk awareness in the return distribution (Lim and Malik 2022) and addressing both aleatoric and epistemic uncertainties (Eriksson and Dimitrakakis 2019; Eriksson et al. 2022).
@@
In this paper, we investigate dynamic risk-awareness strategy with respect to epistemic uncertainty, which reflects the attitude toward exploring unknown environments. Modeling risk-awareness of epistemic uncertainty is particularly relevant for applications such as autonomous driving, where the systems are almost deterministic, and the primary source of uncertainty is the lack of information about the environment. A higher risk-aversion level can avoid the negative consequences of excessive exploration that may lead to unsafe decisions in real-life (pessimism-under-uncertainty). On the other hand, a lower risk-aversion level (including risk-seeking) encourages exploring and collecting more data in unexplored regions (optimism-under-uncertainty). Although standard DRL can learn a policy under different levels of risk-awareness and adapt to various environments that require different degrees of caution, a specific risk-awareness level typically needs to be pre-specified and fixed for each environment prior to deployment. Studies have shown that optimism and pessimism-under-uncertainty settings outperform each other based on the task at hand, highlighting the need for an adaptive risk-awareness strategy that can balance these seemingly opposing perspectives (Moskovitz et al. 2021; Wang, Hu, and Zhang 2024; Wang et al. 2025).
@@
In addition, maintaining a fixed level of risk-awareness throughout the learning period is often practically suboptimal, as the level of risk varies at different stages. An intelligent agent should be able to identify the appropriate level of risk required for effective decision-making at different stages, rather than simply following a fixed risk-aware policy (Schubert et al. 2021; Liu, van Kampen, and De Croon 2023). In the context of RL, problems where uncertainties are expected to diminish over time (e.g., epistemic uncertainty) should be assigned a decreasing risk-awareness level to avoid excessive conservativeness (Eriksson and Dimitrakakis 2019). We illustrate this point further using a Cartpole example.
@@
The Cartpole task was trained through Implicit Quantile Network (IQN) with left-tail Conditional Value-at-Risk (CVaR) as its risk measure on the return (IQN will be explicitly described in Section 3). Over the episodes, the risk parameter $\alpha$ could either be fixed or manually adjusted. Figure 1 shows that “IQN alpha:191” outperforms all other settings, including IQN with fixed risk levels. The intuition is: at early stage of training, set a high risk-awareness level ($\alpha=0.1$) to hedge the uncertain from the unknown environment, then move to a low risk-awareness level ($\alpha=0.9$) to acquire high rewards as more experience from the environment is gained (and the epistemic uncertainty is reduced), and finally set back to a proper value that acts optimally in hindsight. This result highlights the advantage of using varying risk levels and motivates the development of more advanced methods that can automatically adjust the risk level and achieve even better performance.
@@
In this paper, we propose a DRL framework that is capable of automatically adjusting the risk-awareness level on the fly (i.e., online) by leveraging the epistemic uncertainty quantified in each period. The proposed framework does not rely on any pre-specified risk-awareness level and it can be incorporated with a wide range of risk measures based on IQN. Our framework generalizes and improves the existing adaptive risk-awareness level approach. (Moskovitz et al. 2021; Liu, van Kampen, and De Croon 2023; Wang, Hu, and Zhang 2024; Wang et al. 2025). Dynamic risk level selection methods are beneficial for RL algorithms because predetermining a suitable risk-awareness level is difficult, if not impossible, for a new domain task without any prior knowledge. Our experiments demonstrate that the proposed approach, which requires only moderate extensions to standard RL algorithms, achieves superior performance across various classes of practical problems compared to risk-neutral methods, fixed risk-awareness level methods, and existing adaptive risk-awareness level methods.
Related Work
@@
The literature on DRL has significantly expanded in recent years. (Morimura et al. 2010; Bellemare, Dabney, and Munos 2017; Barth-Maron et al. 2018; Dabney et al. 2018a; Yang et al. 2019; Singh, Zhang, and Chen 2020). This approach has been used to enhance exploration in RL (Mavrin et al. 2019) and in risk-aware applications (Bernhard, Pollok, and Knoll 2019). Many different types of risk measures have been used in risk-aware DRL including CVaR (Lim and Malik 2022; Choi et al. 2021; Keramati et al. 2020; Yang et al. 2022), power-law risk measure (Choi et al. 2021), distortion risk measure (Dabney et al. 2018a), entropic risk measure (Hai et al. 2022) and cumulative prospect theory (Prashanth et al. 2016). Research has also been conducted on quantifying the risk of epistemic uncertainty, which demonstrates the attitude towards unexplored areas (Clements et al. 2019; Eriksson and Dimitrakakis 2019; Tang and Agrawal 2018; O’Donoghue 2023), and on establishing comprehensive composite risk, which quantifies the joint effect of aleatory and epistemic risk during the learning process (Eriksson et al. 2022; Cho et al. 2023).
@@
The above literature focuses on fixed risk-awareness levels. Recently, there has been growing evidence supporting the adaptive risk-awareness algorithms, such as risk scheduling approaches that explore risk levels and optimistic behaviors from a risk perspective (Oh, Kim, and Yun 2022). These methods are not automatic, however, risk levels decay linearly from high to specific low levels during the learning process. In addition, frameworks for value estimation, such as Tactical Optimism and Pessimism (TOP) and Quantile Option Architecture (QUOTA) (Moskovitz et al. 2021; Wang, Hu, and Zhang 2024; Wang et al. 2025; Zhang and Yao 2019), have been developed to balance the degree of optimistic and pessimistic uncertainty quantification. Many of these frameworks formulate the problem as a multi-armed bandit problem and utilize the Exponentially Weighted Average Forecasting (EWAF) algorithm. To implement this method, a pre-specified degree level should be set based on expert recommendations. A similar EWAF algorithm is used for adaptive risk tendency (ART) of CVaR in an IQN framework for a Nano Drone navigation application (Liu, van Kampen, and De Croon 2023), and it achieves superior reward accumulation and efficient computational performance. The drawback of this method appears during evaluation in environments with high obstacle density, possibly because it relies solely on the current observation without utilizing any memory. Besides, EWAF is not intuitive , and tuning the learning rate for weight forecasting still requires human effort. In another work (Schubert et al. 2021), risk levels are adjusted based on a random network distillation (RND) error, which provides an estimate of how frequently a state has been visited during training. While the performance heavily depends on the original RND architecture, it is possible that a more conservative risk setting could hinder exploration.
Preliminaries
@@
Consider a Markov Decision Process (MDP) (Puterman 2014) defined as a tuple $(\mathbb{S},\mathbb{A},R,P,\gamma)$, where $\mathbb{S}$ is the state space, $\mathbb{A}$ the action space, $R$ is the (state and action-dependent) reward function, $P(\cdot|s,\,a)$ is the transition probability associated with state-action pair (s,a)(s,\,a), and $\gamma\in(0,1]$ is the discounted factor. A policy $\pi(\cdot|s)$ maps the state $s$ to a probability distribution over actions.
DRL (Bellemare, Dabney, and Munos 2017) introduces a new perspective other than the standard RL setting, by learning the entire return distribution of the action-value function Q(s,a)Q(s,\,a) directly instead of estimating only the expected return Q(s,a)Q(s,\,a). This approach allows the agent to capture more detailed information beyond a single expectation value. Given a policy $\pi$, the distributional Bellman equation is defined as
@@
for $s^{\prime}\sim P(\cdot\mid s,\,a)$. It is proved by Bellemare, Dabney, and Munos (2017) that the distributional Bellman operator is a contraction in a maximal form of the Wasserstein metric. There are approaches (Bellemare, Dabney, and Munos 2017; Dabney et al. 2018b, a; Yang et al. 2019; Rowland et al. 2019) to estimate the return distribution, most of which are based on Quantile Regression (Koenker and Bassett Jr 1978).
The Implicit Quantile Network (IQN) proposed by Dabney et al. (2018a) provides an effective way to learn an implicit representation of the return distribution, yielding a powerful function approximator for a Deep $Q$-Network agent. IQN can re-parameterize samples from a based distribution to the respective quantile value of a target distribution, which produces risk-aware policies. Let $\beta:[0,\,1]\to[0,\,1]$ be a distortion function, and choosing $\beta$ as the identity function corresponds to risk-neutrality. Let $Z_{q}(s,\,a)$ denote the $q$-quantile function (inverse of cumulative density function) of the return distribution, then the distorted expectation of Z(s,a)Z(s,\,a) under $\beta$ and its risk-aware greedy policy are denoted as
@@
The proposed adaptive risk-awareness approach treats the selection of an exploration strategy (risk awareness level) as a multi-armed bandit problem (Moskovitz et al. 2021; Wang, Hu, and Zhang 2024; Wang et al. 2025), where each bandit arm corresponds to a specific choice of distortion function $\beta$. Define a discrete set of $D$ distortion functions $\mathcal{D}=\{\beta^{1},...,\beta^{D}\}$ recommended by experts. After sampling $d_{n}\in\{1,...,D\}$ at epsiode $n$, we form a belief distribution $p_{n}\in\Delta_{\mathcal{D}}$ ($\Delta_{\mathcal{D}}$ is the probability simplex on $\mathcal{D}$) of the form $p_{n}(d)\propto\exp(w_{n}(d))$ for all $d\in\mathcal{D}$. The distribution is updated using the Exponentially Weighted Average Forecasting (EWAF) algorithm based on a received feedback signal in each episode. After receiving a feedback signal $g_{n}:=R_{n}-R_{n-1}$ where $R_{n}$ is the cumulative reward obtained in episode $n$, the weight $w_{n}$ is updated by:
$w_{n+1}(d)=\begin{cases}w_{n}(d)+\eta\cdot(g_{n}/p(d)),&d=d_{n},\\ w_{n}(d),&\text{otherwise},\end{cases}$
@@
Another variants of constructing the feedback signal is the right-truncated variance difference of reward distributions between two successive periods (time-steps) (Liu, van Kampen, and De Croon 2023), where $g_{t}:=\text{RTV}_{t}-\text{RTV}_{t-1}$ and
$\text{RTV}_{t}:=\frac{2}{N}\sum_{i=1}^{N/2}(F^{-1}_{Z_{t}}(q_{i})-F^{-1}_{Z_{t}}(q_{N/2})^{2},$
@@
Methodology
@@ A Generalized Non-convex Learning Perspective
In aforementioned EWAF approach, the major shortcoming is its explainability. The theoretical foundation / objective of the approach in updating $\alpha$ is unclear. The most similar framework to the approach is online optimization on simplex (Bubeck 2011), where the signals must be an explicit loss function in the decision variable, rather than simple numerical outputs.
There are also several other shortcomings. In this approach, epistemic uncertainty is not disentangled from the reward distribution, which means that the agent cannot specify the risk-awareness level towards exploration and further leads to inaccurate adaptation. Besides, it relies on truncated variance or cumulative reward, rather than leveraging the full information from the epistemic uncertainty distribution, to construct the feedback signal. Furthermore, this approach requires a predefined discrete finite set of distortion functions (e.g., only two functions are used in (Liu, van Kampen, and De Croon 2023)).
@@
where $X_{t}(s,\,a)$ represents the epistemic uncertainty associated with the state-action pair (s,a)(s,\,a). Here we define the parametric risk measure on the epistemic uncertainty: $\rho_{\alpha}:\mathcal{L}\rightarrow\mathbb{R}$, where $\mathcal{L}$ represents the space of bounded random variables, $\alpha\in\mathcal{A}$ is a parameter that controls the level of risk-awareness, and $\mathcal{A}\subset\mathbb{R}$ is a compact set containing all possible risk parameters. Explicitly, $\alpha(s,\,a)\in\mathcal{A}$ denotes the risk parameter chosen corresponding to the state-action pair (s,a)(s,\,a).
The risk measure $\rho_{\alpha}(\cdot)$ is Lipschitz continuous but not necessarily convex with respect to $\alpha$. Distortion risk measures (Wang 1996; Dabney et al. 2018b) fall under our framework and are well-suited for the IQN framework. Specifically, for $X\in\mathcal{L}$, we have
@@
Example 2 ((Acerbi and Tasche 2002)).
For the right-tail CVaR, $\text{CVaR}_{\alpha}(X)=\mathbb{E}[X\mid X\geq F^{-1}(1-\alpha)]$, where $F^{-1}$ denotes the inverse CDF of $X$. The corresponding distortion function is $g_{\alpha}(x)=\max\left\{-x/\alpha+1,\,0\right\}$, and $\mathcal{A}=[\alpha_{\min},\,1]$ for some $\alpha_{\min}\in(0,1]$.
Example 3 ((Dhaene et al. 2012)).
@@
    Our feedback signal generalizes the existing approach by utilizing the full information of the epistemic uncertainty distribution. Also, the feedback signal is defined as an explicit function of the risk parameter, so the risk adaption problem can be formally written as an online learning problem with clear objective and decision foundation.
@@
    Our approach provides higher flexibility, as the risk levels associated with different state-action pairs can be updated independently and simultaneously. Furthermore, the adaptation is performed in every period rather than episodically, as in (Moskovitz et al. 2021; Wang, Hu, and Zhang 2024; Wang et al. 2025). These will be validated later by experiments, demonstrating improved training performance during early periods and episodes.
@@
    The risk measure $\rho_{\alpha}(\cdot)$ focuses on the right $\alpha$-level quantile of a distribution, capturing the risk of left-truncated epistemic uncertainty (reflecting upper-tail variability), which makes it comparable with the setting of reward distribution RTV in (Liu, van Kampen, and De Croon 2023). It is inspired by setting a contrary to (Mavrin et al. 2019), which argues that the upper-tail variability of reward distribution reflects optimism of exploration in the face of uncertainty, particularly when the distribution is asymmetric.
@@
which can be interpreted as finding a state-action-dependent risk parameter that minimizes the Total Variation (Chambolle 2004; Huang, Ng, and Wen 2008; Li et al. 2013) (i.e., the sum of feedback signals) of the epistemic uncertainty risk.
@@
Building on the principles of Bayesian inference (MacKay 2003), the ensemble network is a straightforward approach for modeling epistemic uncertainty and is commonly used in prior work (Dimitrakakis 2006; Osband et al. 2016; Agarwal, Schuurmans, and Norouzi 2020). Recall that in the regular RL framework, a neural network $\theta$ is used to estimate the $Q$-function $Q^{\theta}(s,a)$, where the distribution of $\theta$ reflects the current epistemic uncertainty during training (Eriksson et al. 2022). While modeling the distribution of $\theta$ is challenging, its corresponding network outputs $Q^{\theta}(s,\,a)$ can be directly observed.

### v5

# DRL-ORA: Distributional Reinforcement Learning with Online Epistemic Risk Adaptation

@@ Abstract
One of the main challenges in reinforcement learning (RL) is that the agent has to make decisions that would influence the future performance without having complete knowledge of the environment. Dynamically adjusting the level of epistemic risk during the learning process can help to achieve reliable policies in safety-critical settings with better efficiency. In this work, we propose a new framework, Distributional RL with Online Epistemic Risk Adaptation (DRL-ORA). This framework quantifies both epistemic and implicit aleatory uncertainties in a unified manner and dynamically adjusts the epistemic risk levels by solving a total variation minimization problem online. The framework generalizes the existing variants of risk adaptation approaches with better explainability and flexibility. The selection of risk levels is performed efficiently via a Follow-The-Leader-type algorithm.We show that DRL-ORA outperforms existing methods that rely on fixed risk levels or manually designed risk level adaptation in multiple classes of tasks.
1 Introduction
@@
Reinforcement learning (RL) algorithms have shown great success in games and simulated environments [mnih2015human], and have drawn interest for real-world and industrial applications [mahmood2018benchmarking, zhang2020cautious]. Taking a step further, distributional reinforcement learning (DRL) [bellemare2017distributional] considers the intrinsic distribution of the returns rather than their expectation, as considered in standard RL settings. This approach allows agents to quantify the risk of their policies and enables the development of risk-aware policies by considering the distributional information of their performance [dabney2018distributional]. Quantifying both aleatory and epistemic uncertainties of RL crucial for developing risk-aware policies. The former expresses the inherent randomness in the problem, and the latter represents a lack of knowledge about the environment [mavrin2019distributional, clements2019estimating, charpentier2022disentangling]. However, most existing research has focused on risk-aware RL with a fixed risk level, such as investigating risk awareness in the return distribution [lim2022distributional] and addressing both aleatoric and epistemic uncertainties [eriksson2019epistemic, eriksson2022sentinel].
@@
In this paper, we investigate dynamic risk-awareness strategy with respect to epistemic uncertainty, which reflects the attitude toward exploring unknown environments. Modeling risk-awareness of epistemic uncertainty is particularly relevant for applications such as autonomous driving, where the systems are almost deterministic, and the primary source of uncertainty is the lack of information about the environment. A higher risk-aversion level can avoid the negative consequences of excessive exploration that may lead to unsafe decisions in real-life (pessimism-under-uncertainty). On the other hand, a lower risk-aversion level (including risk-seeking) encourages exploring and collecting more data in unexplored regions (optimism-under-uncertainty). Although standard DRL can learn a policy under different levels of risk-awareness and adapt to various environments that require different degrees of caution, a specific risk-awareness level typically needs to be pre-specified and fixed for each environment prior to deployment. Studies have shown that optimism and pessimism-under-uncertainty settings outperform each other dependent on the task at hand, highlighting the need for an adaptive risk-awareness strategy that can balance these seemingly opposing perspectives [moskovitz2021tactical, wang2024improving, wang2025adaptive].
@@
In addition, maintaining a fixed level of risk-awareness throughout the learning period is often practically suboptimal, as the level of risk concerned varies at different stages. An intelligent agent should be able to identify the appropriate level of risk required for effective decision-making at different stages, rather than simply following a fixed risk-aware policy [schubert2021automatic, liu2022adaptive]. In the context of RL, problems where uncertainties are expected to diminish over time (e.g., epistemic uncertainty) should be assigned a decreasing risk-awareness level to avoid excessive conservativeness [eriksson2019epistemic]. We illustrate this point further using a CartPole example.
@@
The CartPole task was trained through Implicit Quantile Network (IQN) with left-tail Conditional Value-at-Risk (CVaR) at level $\alpha\in(0,\,1]$ as its risk measure on the return (IQN will be explicitly described in Section 3). Over the episodes, the risk parameter $\alpha$ could either be fixed or manually adjusted. Figure 1 shows that “IQN alpha:191” outperforms all other settings, including IQN with fixed risk levels. The intuition is: at early stage of training, set a high risk-awareness level ($\alpha=0.1$) to hedge the uncertain from the unknown environment, then move to a low risk-awareness level ($\alpha=0.9$) to acquire high rewards as more experience from the environment is gained (and the epistemic uncertainty is reduced), and finally set back to a value in [0, 1][0,\,1] that acts optimally in hindsight. This result highlights the advantage of using varying risk levels and motivates the development of more advanced methods that can automatically adjust the risk level and achieve even better performance.
@@
In this paper, we propose a DRL framework that is capable of automatically adjusting the risk-awareness level on the fly (i.e., online) by the epistemic uncertainty quantification in each period. The proposed framework does not rely on any pre-specified risk-awareness level and it can be incorporated with a wide range of risk measures based on IQN. Our framework generalizes and improves the existing adaptive risk-awareness level approach [moskovitz2021tactical, liu2022adaptive, wang2024improving, wang2025adaptive]. Dynamic risk level selection methods are beneficial for RL algorithms because predetermining a suitable risk-awareness level is difficult, if not impossible, for a new domain task without any prior knowledge. Our experiments demonstrate that the proposed approach, which requires only moderate extensions to standard RL algorithms, achieves superior performance across various classes of practical problems compared to risk-neutral methods, fixed risk-awareness level methods, and existing adaptive risk-awareness level methods.
2 Related Work
@@
The literature on DRL has significantly expanded in recent years [morimura2010parametric, bellemare2017distributional, barth2018distributed, dabney2018implicit, yang2019fully, singh2020improving]. This approach has been used to enhance exploration in RL [mavrin2019distributional] and in risk-aware applications [bernhard2019addressing]. Many different types of risk measures have been used in risk-aware DRL including CVaR [lim2022distributional, choi2021risk, keramati2020being, yang2022safety], power-law risk measure [choi2021risk], distortion risk measure [dabney2018implicit], entropic risk measure [hai2022rasr] and cumulative prospect theory [prashanth2016cumulative]. Research has also been conducted on quantifying the risk of epistemic uncertainty, which demonstrates the attitude towards unexplored areas [clements2019estimating, eriksson2019epistemic, tang2018exploration, o2023efficient], and on establishing comprehensive composite risk, which quantifies the joint effect of aleatory and epistemic risk during the learning process [eriksson2022sentinel, cho2023pitfall].
@@
The above literature focuses on fixed risk-awareness levels. Recently, there has been growing evidence supporting the adaptive risk-awareness algorithms, such as risk scheduling approaches that explore risk levels and optimistic behaviors from a risk perspective [oh2022risk]. These methods are not situation-dependent, however, risk levels decay linearly from high to specific low levels during the learning process. In addition, frameworks for value estimation, such as Tactical Optimism and Pessimism (TOP) and Quantile Option Architecture (QUOTA) [moskovitz2021tactical, wang2024improving, wang2025adaptive, zhang2019quota], have been developed to balance the degree of optimistic and pessimistic uncertainty quantification. Many of these frameworks formulate the problem as a multi-armed bandit problem and utilize the Exponentially Weighted Average Forecasting (EWAF) algorithm. To implement this method, a pre-specified degree level should be set based on expert recommendations. A similar EWAF algorithm is used for adaptive risk tendency (ART) of CVaR in an IQN framework for a Nano Drone navigation application [liu2022adaptive], and it achieves superior reward accumulation and efficient computational performance. The drawback of this method appears during evaluation in environments with high obstacle density, possibly because it relies solely on the current observation without utilizing any memory. Besides, EWAF is not intuitive , and tuning the learning rate for weight forecasting still requires extra effort. In another work [schubert2021automatic], risk levels are adjusted based on a random network distillation (RND) error, which provides an estimate of how frequently a state has been visited during training. While the performance heavily depends on the original RND architecture, it is possible that a more conservative risk setting could hinder exploration.
3 Preliminaries
@@
Consider a Markov Decision Process (MDP) [puterman2014markov] defined as a tuple $(\mathbb{S},\mathbb{A},R,P,\gamma)$, where $\mathbb{S}$ is the state space, $\mathbb{A}$ the action space, $R$ is the (state and action-dependent) reward function, $P(\cdot|s,\,a)$ is the transition probability associated with state-action pair (s,a)(s,\,a), and $\gamma\in(0,1]$ is the discounted factor. A policy $\pi(\cdot|s)$ maps the state $s$ to a probability distribution over actions.
DRL [bellemare2017distributional] introduces a new perspective other than the standard RL setting, by learning the entire return distribution of the action-value function Q(s,a)Q(s,\,a) directly instead of estimating only the expected return Q(s,a)Q(s,\,a). This approach allows the agent to capture more detailed information beyond a single expectation value. Given a policy $\pi$, the distributional Bellman equation is defined as
@@
for $s^{\prime}\sim P(\cdot\mid s,\,a)$. It is proved by bellemare2017distributional that the distributional Bellman operator is a contraction in a maximal form of the Wasserstein metric. There are approaches [bellemare2017distributional, dabney2018distributional, dabney2018implicit, yang2019fully, rowland2019statistics] to estimate the return distribution, most of which are based on Quantile Regression [koenker1978regression].
The Implicit Quantile Network (IQN) proposed by dabney2018implicit provides an effective way to learn an implicit representation of the return distribution, yielding a powerful function approximator for a Deep $Q$-Network agent. IQN can re-parameterize samples from a based distribution to the respective quantile value of a target distribution, which produces risk-aware policies. Let $\beta:[0,\,1]\to[0,\,1]$ be a distortion function, and choosing $\beta$ as the identity function corresponds to risk-neutrality. Let $Z_{q}(s,\,a)$ denote the $q$-quantile function (inverse of cumulative density function) of the return distribution, then the distorted expectation of Z(s,a)Z(s,\,a) under $\beta$ and its risk-aware greedy policy are denoted as
@@
The proposed adaptive risk-awareness approach treats the selection of an exploration strategy (risk awareness level) as a multi-armed bandit problem [moskovitz2021tactical, wang2024improving, wang2025adaptive], where each bandit arm corresponds to a specific choice of distortion function $\beta$. Define a discrete set of $D$ distortion functions $\mathcal{D}=\{\beta^{1},...,\beta^{D}\}$ recommended by experts. After sampling $d_{n}\in\{1,...,D\}$ at epsiode $n$, we form a belief distribution $p_{n}\in\Delta_{\mathcal{D}}$ ($\Delta_{\mathcal{D}}$ is the probability simplex on $\mathcal{D}$) of the form $p_{n}(d)\propto\exp(w_{n}(d))$ for all $d\in\mathcal{D}$. The distribution is updated using the Exponentially Weighted Average Forecasting (EWAF) algorithm based on a received feedback signal in each episode. After receiving a feedback signal $g_{n}:=R_{n}-R_{n-1}$ where $R_{n}$ is the cumulative reward obtained in episode $n$, the weight $w_{n}$ is updated by:
$w_{n+1}(d)=\begin{cases}w_{n}(d)+\eta\cdot(g_{n}/p_{n}(d)),&d=d_{n},\\ w_{n}(d),&\text{otherwise},\end{cases}$
@@
Another variants of constructing the feedback signal is the right-truncated variance difference of reward distributions between two successive periods (time-steps) [liu2022adaptive], where $g_{t}:=\text{RTV}_{t}-\text{RTV}_{t-1}$ and
$\text{RTV}_{t}:=\frac{2}{N}\sum_{i=1}^{N/2}(F^{-1}_{Z_{t}}(q_{i})-F^{-1}_{Z_{t}}(q_{N/2}))^{2},$
@@
4 Methodology: DRL-ORA
We propose DRL-ORA, the first distributional reinforcement learning framework capable of online, state-action-level adaptation of epistemic risk attitudes without requiring pre-specified risk levels or manual scheduling. At its core, DRL-ORA introduces a unified uncertainty quantification scheme that explicitly disentangles epistemic from aleatory uncertainty via ensemble networks, and formulates risk adaptation as a non-convex online learning problem with a well-defined regret objective. This marks a fundamental departure from prior approaches—such as EWAF-based bandit selection—which lack explainability, rely on predefined discrete risk sets, and fail to leverage full distributional information. In contrast, DRL-ORA enables fine-grained, per-transition risk adaptation, grounded in a total variation minimization criterion that is both theoretically tractable and computationally efficient. The framework is agnostic to the choice of distortion risk measures, supports both CVaR and quantile-based variants, and establishes a novel connection to satisficing measures in decision theory—revealing that the offline oracle in our formulation corresponds to a quasi-concave satisficing objective. Together, these innovations yield a flexible, interpretable, and high-performing risk adaptation paradigm that consistently outperforms fixed-risk and existing adaptive baselines across diverse tasks.
@@ A Generalized Non-convex Learning Perspective
In the existing EWAF approach, the major shortcoming is its explainability. The theoretical foundation/objective of the approach in updating the risk tendency is unclear. The most similar framework to the approach is online optimization on simplex [bubeck2011introduction], where the signals must be an explicit loss function in the decision variable, rather than simple numerical outputs. There are also several other shortcomings. In this approach, epistemic uncertainty is not disentangled from the reward distribution, which means that the agent cannot specify the risk-awareness level towards exploration as it will further lead to inaccurate adaptation. Besides, it relies on truncated variance or cumulative reward, rather than leveraging the full information from the epistemic uncertainty distribution, to construct the feedback signal. Furthermore, this approach requires a predefined discrete finite set of distortion functions (e.g., only two functions are used in [liu2022adaptive]).
@@
where $X_{t}(s,\,a)$ represents the epistemic uncertainty associated with the state-action pair (s,a)(s,\,a) at period $t$. Here we define the parametric risk measure on the epistemic uncertainty: $\rho_{\alpha}:\mathcal{L}\rightarrow\mathbb{R}$, where $\mathcal{L}$ represents the space of bounded random variables, $\alpha\in\mathcal{A}$ is a parameter that controls the level of risk-awareness, and $\mathcal{A}\subset\mathbb{R}$ is a compact set containing all possible risk parameters. Explicitly, $\alpha(s,\,a)\in\mathcal{A}$ denotes the risk parameter chosen corresponding to the state-action pair (s,a)(s,\,a).
The risk measure $\rho_{\alpha}(\cdot)$ is Lipschitz continuous but not necessarily convex with respect to $\alpha$. Distortion risk measures [wang1996premium, dabney2018distributional] fall under our framework and are well-suited for the IQN framework. Specifically, for $X\in\mathcal{L}$, we have
@@
Example 2 ([acerbi2002coherence]).
For the right-tail CVaR, $\text{CVaR}_{\alpha}(X)=\mathbb{E}[X|X\geq F^{-1}(1-\alpha)]$, where $F^{-1}$ denotes the inverse CDF of $X$. The corresponding distortion function is $g_{\alpha}(x)=\max\left\{-x/\alpha+1,\,0\right\}$, and $\mathcal{A}=[\alpha_{\min},\,1]$ for some $\alpha_{\min}\in(0,1]$.
Example 3 ([dhaene2012remarks]).
@@
    Our feedback signal generalizes the existing approach by utilizing the full information of the epistemic uncertainty distribution. Also, the feedback signal is defined as an explicit function of the risk parameter, so the risk adaptation problem can be formally written as an online learning problem with clear objective and decision foundation.
@@
    Our approach provides higher flexibility, as the risk levels associated with different state-action pairs can be updated independently and simultaneously. Furthermore, the adaptation is performed in every period/transition rather than episodically, as in [moskovitz2021tactical, wang2024improving, wang2025adaptive]. These will be validated later by experiments, demonstrating improved training performance during early periods and episodes.
@@
    The risk measure $\rho_{\alpha}(\cdot)$ focuses on the right $\alpha$-level quantile of a distribution, capturing the risk of left-truncated epistemic uncertainty (reflecting upper-tail variability), which makes it comparable with the setting of reward distribution RTV in [liu2022adaptive]. It is inspired by setting a contrary to [mavrin2019distributional], which argues that the upper-tail variability of reward distribution reflects optimism of exploration in the face of uncertainty, particularly when the distribution is asymmetric.
@@
which can be interpreted as finding a state-action-dependent risk parameter that minimizes the Total Variation [chambolle2004algorithm, huang2008fast, li2013efficient] (i.e., the sum of feedback signals) of the epistemic uncertainty risk.
@@
Building on the principles of Bayesian inference [mackay2003information], the ensemble network is a straightforward approach for modeling epistemic uncertainty and is commonly used in prior work [dimitrakakis2006nearly, osband2016deep, agarwal2020optimistic]. Recall that in the regular RL framework, a neural network $\theta$ is used to estimate the $Q$-function $Q^{\theta}(s,a)$, where the distribution of $\theta$ reflects the current epistemic uncertainty during training [eriksson2022sentinel]. While modeling the distribution of $\theta$ is challenging, its corresponding network outputs $Q^{\theta}(s,\,a)$ can be directly observed.